OpenStudy (anonymous):

What is the limit as x approaches negative infinity for the function (sqrt(x^2 + x)) - x ?

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

negative infinity...

OpenStudy (anonymous):

nope! its 1/2 as told by him :)

OpenStudy (anonymous):

chalo yeh to aata hai

OpenStudy (anonymous):

as x goes to MINUS infinity?

OpenStudy (anonymous):

yes maam

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it doesn't matter if it goes to - or + infinity here :)

OpenStudy (anonymous):

really? i believe you but i don't see it. first term goes to infinity. second term also goes to infinity. why 1/2?

OpenStudy (anonymous):

maam? oh! u never told ur a prof. sat !

OpenStudy (anonymous):

more to it than meets the eye

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

no praveen m born chivalric..i can call neone maam

OpenStudy (anonymous):

ah! its cause it not addtion rather subtraction...u need to rationalise and then proceed to pull out an x .

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

ah! i'm going to crash , now sat is correct!

OpenStudy (anonymous):

whole section in Amit Agarwal's "Differential Calculus" on these types of qstns

OpenStudy (anonymous):

pellet! i've not slept for 2 days...i'm flipping out :) okay guys night ! and sat-u stand right :)

OpenStudy (watchmath):

1/2 is wrong because as x--> -infinity \(\sqrt{x^2-x}=-x\sqrt{1-(1/x)}\)

OpenStudy (anonymous):

but it is not \[\infty - \infty\] is it \[\infty + \infty\]

OpenStudy (anonymous):

yea as i told u sat ur right :)

OpenStudy (anonymous):

i still stand by 1/2..how is sat right?

OpenStudy (watchmath):

it is true if it goes to +infinity

OpenStudy (anonymous):

1/2 is right only wen it tends to + inf. and not -vee inf. just let the value tend to - inf. u'll see it tending to +inf :) salaute sat \./

OpenStudy (anonymous):

ema123: answer is finite limit doesn't exist .

OpenStudy (anonymous):

ur left with \[1/ [\sqrt{1 + 1/x} + 1]\]

OpenStudy (anonymous):

now if u put in -inf into it the 1/x tends to zero doesnt it?

OpenStudy (watchmath):

I told you above when you pull the x outside the radical it become -x since your x is negative.

OpenStudy (anonymous):

If you multiply times the conjugate then by 1/x / 1/x you get 1/(sqrt(x^2+x)+1) as x-> infit you get 1/2. To find x -> -infit you can instead do as x-> infit of f(-x) which you still get 1/2 because 1/-x still goes to 0

OpenStudy (watchmath):

Maybe you believe wolframalpha him? http://www.wolframalpha.com/input/?i=lim_ {x%5Cto+-%5Cinfty}+%28%5Csqrt{x^2%2Bx}-x%29

OpenStudy (watchmath):

http://www.wolframalpha.com/input/?i=lim_ {x%5Cto+-%5Cinfty}+%28%5Csqrt{x^2%2Bx}-x%29

OpenStudy (watchmath):

sorry the link is not working :(

OpenStudy (anonymous):

see himanshi: u can take these rationalisation only wen there is ambiguity in the value, here its direct saying it tends to inf.

OpenStudy (anonymous):

i have a solved exaple in the book

OpenStudy (anonymous):

you don't need to continues unless you have something "indeterminate" if you get infinity by inspection that is the answer.

OpenStudy (anonymous):

take a pic and post it.we'll tell if the author is wrong.

OpenStudy (anonymous):

but here u get infty - infty so u need to rationalize as its an indeterminate form

OpenStudy (anonymous):

thats the mistake ur makin....the form is no more inf-inf, its inf+inf since x->-inf... -(-) is +

OpenStudy (anonymous):

no because the limit is as x goes to \[-\infty\] not \[\infty\]

OpenStudy (anonymous):

get it now

OpenStudy (anonymous):

so as x goes to \[-\infty\] both the first and second term go to \[\infty\] done

OpenStudy (anonymous):

oh yes right..now i c it